OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find a if b = 6 and c = 10. a. a = 8 c. a = 64 b. a = 4 d. a = 16

OpenStudy (anonymous):

use pythagorean theory to solve a^2+b^2=c^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what did you get for answer

OpenStudy (anonymous):

64

OpenStudy (anonymous):

great thats what I got too

OpenStudy (anonymous):

okay so then c..... thanks.

OpenStudy (anonymous):

the only thing that confuses me and it might be the answer too is the answer a. a=8 because a^2 =64 so I am not sure if they want you to use that as b and c are listed as 6 and 10 which would be before they are squared maybe someone else can check this

OpenStudy (anonymous):

8^2=64

OpenStudy (anonymous):

yeah thats what i figured it was

OpenStudy (anonymous):

@Directrix what would you use as the answer to this figured out the length was 64 but would you use the answer a. a=8 which would be 64 after you squared it

OpenStudy (anonymous):

oh sorry before you squared it

Directrix (directrix):

The missing side is just 8. Is that what you are asking?

OpenStudy (anonymous):

yes thanks

Directrix (directrix):

Okay, then.